OpenStudy (anonymous):

Can someone help please!! What are the approximate solutions of 4x^2+3=-12x to the nearest hundredth? Super confused 😁😁😁

OpenStudy (here_to_help15):

To solve a quadratic equation, put the trinomial on one side and make it equal to 0. Then factor it and make each set of parentheses equal to 0. Then solve both equations to get your solution. 4x² + 3 = -12x +12x = +12x 4x² + 12x + 3 = 0 x = -0.28 , -2.73

OpenStudy (anonymous):

Thank you